OpenStudy (anonymous):

Add the Algebraic Fractions and Reduce to lowest terms: 4/(a+3)+2/(3a) this is the choices 12a+2/3a(a+3) 14a+6/a+3 12a+2/3a 14a+6/3a(a+3)

OpenStudy (anonymous):

Find the LCM of the fractions and you get the answer! 4/(a+3)+2/(3a) = (4*3a+2*(a+3))/3a(a+3) =(14a+6)/3a(a+3)